# Bulk viscosity of nuclear matter with pions in the neutrino-trapped regime

### v1

@@ IV.2 Particle dynamics during an oscillation
The evolution of the particle fractions $x_{p}(t)$, $x_{\mu}(t)$, and $x_{\pi}(t)$ (more precisely, the quantity $(x_{i}(t)-x_{i}^{0})/x_{i}^{0}$ where $x_{i}^{0}$ is the beta-equilibrium value of the particle fraction of species $i$) are plotted in Fig. 10 at different temperatures and densities, with and without pions in the EoS. In studying these plots, both the height and the phase of the sinusoidal variations in the particle fractions are informative. If the equilibration of a species is very slow, then its deviation from its beta equilibrium value will be very small, proportional to $\gamma_{i}/\omega$. To the extent that the particle fraction oscillates, it is out of phase with the density oscillation. As the equilibration rate increases (for example, with rising temperature), then the particle fraction deviates more from its beta equilibrium value and its oscillation becomes more in phase with the density oscillation, becoming completely in phase in the limit of infinitely fast equilibration. During a period, the maximum amount of deviation of the particle fraction from equilibrium becomes independent of the equilibration rate as the equilibration rate goes to infinity. That is, how far a particle fraction deviates from equilibrium depends just on the properties of the EoS (the susceptibilities) and the density oscillation amplitude in this limit. We see in Fig. 10 that the particle fraction oscillations become increasingly in phase with the density oscillation as temperature increases.
@@ V Conclusions
Our main finding is that thermal pions can significantly alter the density and temperature dependence of the bulk viscosity of hot dense matter. The pion population rapidly increases with density and temperature, and so the bulk viscosity deviates most strongly from the pionless case at higher densities and temperatures. The addition of pions to the EoS modifies the existing nuclear matter susceptibilities to a relatively small extent (Figs. 3 and 4) but also introduces new susceptibilities that do not exist without pions. These new susceptibilities have a significant effect on the bulk viscosity. In addition, pion reaction rates have two different effects. The (essentially) infinitely fast strong interaction $n+n\rightarrow n+p+\pi^{-}$ does not produce a new resonant peak in the bulk viscosity in the temperature range studied, but does (as one can see from Fig. 5) modify the maximum value of the bulk viscosity and can move or even eliminate or introduce partial conformal points in the bulk viscosity at fixed density, $\zeta(T)$. The slower pion decays ($\pi^{-}\rightarrow\mu^{-}+\bar{\nu}_{\mu}$) in some conditions can overtake the existing nucleonic or leptonic equilibration processes and can dominate beta equilibration.
@@ Acknowledgements.
SPH thanks Mark Alford for collaboration in the early stages of the project, and Albino Perego and Eleonora Loffredo for information about the conserved lepton fractions in neutron star mergers. The work of SPH was supported by the U.S. Department of Energy grant DE-FG02-00ER41132 as well as the National Science Foundation grants PHY-1430152 (JINA Center for the Evolution of the Elements) and PHY 21-16686. BF is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under contracts DE-AC02-06CH11357, by the 2020 DOE Early Career Award number ANL PRJ1008597, by the NUCLEI SciDAC program, and Argonne LDRD awards. The work of SR was supported by the U.S. Department of Energy grant DE-FG02- 00ER41132. SR also thanks the members of the N3AS Physics Frontier Center, funded by the NSF Grant No. PHY-2020275 for useful conversations.
@@ Appendix A Bulk viscosity of matter with net muon lepton fraction $Y_{L\mu}$
When $Y_{Le}=Y_{L\mu}=0.05$, the electron and muon populations are roughly equal. At $n_{0}$, the proton fraction is 10% instead of around 5% (in the $Y_{L\mu}=0$ case). When pions are added to the system, their population exponentially rises with temperature, becoming appreciable at $T\approx 20\text{ MeV}$. When the pion population becomes appreciable, the proton fraction is raised and the lepton fractions are slightly lowered. At $1.6n_{0}$, the pions content becomes appreciable at very low temperatures, just a couple MeV. Thus the proton fraction is larger than without pions for the basically the entire temperature range plotted.
@@
Figure 12: Susceptilities of the $A$ type from the EoS described in Sec. II plotted in merger conditions (top panels) and in supernovae conditions (bottom panels), but with equal conserved lepton fractions. The dashed lines correspond to an EoS without pions, while the solid lines indicate an EoS with pions. $A_{3}$ has no counterpart in a system without pions.
Figure 13: Susceptilities of the $B$, $C$, and $D$ types from the EoS described in Sec. II plotted in merger conditions (top panels) and in supernovae conditions (bottom panels), but with equal conserved lepton fractions. The dashed lines correspond to an EoS without pions, while the solid lines indicate an EoS with pions. $D_{3}$ has no counterpart in a system without pions.
@@
The equilibration rates $\gamma_{i}$ are plotted in Fig. 14. In merger conditions, the rates are much faster when $Y_{L\mu}=0.05$ than when $Y_{L\mu}=0$. Adding pions to the EoS in these equal lepton fraction conditions does not modify the beta equilibration rates much, except adding the $n+n\leftrightarrow n+p+\pi^{-}$ equilibration channel which (we assume) is infinitely fast. In supernovae conditions, the rates are hardly modified at all, because the pion content is very low.
@@ B.1 Direct Urca (with electrons)
The phase space integral reduces from 12 to 8 dimensions after integrating over the delta functions. Then a spherical coordinate system can be chosen in such a way to render three additional angular integrals trivial, leaving a 5-dimensional integral. Two more integrals can be done analytically, leaving a 3-dimensional integral to do numerically. This integral has been done with the full matrix element (Eq. 35) with $g_{A}=1$ in [4], with the non-relativistic matrix element 36 in [5], and in the neutrino-transparent case in [50, 49].
@@ B.2 Pion decay to electron
The rate of the process $\pi^{-}\rightarrow\mu^{-}+\bar{\nu}_{\mu}$ can be obtained by taking the rate integral 38 with matrix element 39 and replacing the electron mass and chemical potential with those of the muon, as well as replacing the electron neutrino chemical potential with the muon neutrino chemical potential.
@@ Appendix C Maxwell Relations
From this 1st law of thermodynamics, we can derive six Maxwell relations (if we only consider derivatives with respect to $n_{B}$ or $x_{i}$) [67]. The three that are relevant are written in Eqs. 21a, 21b, and 21c.
@@ Appendix F Bulk viscosity $\lambda_{3}\rightarrow\infty$
$+\lambda_{6}\bigg{[}A_{2}(B_{1}-B_{2})+A_{1}C_{2}-\dfrac{A_{3}B_{2}(B_{1}-B_{2}+C_{2})}{B_{2}+D_{3}}\bigg{]}^{2}\Big{\}}.$
@@
$+2(B_{1}-B_{2})^{2}(\lambda_{1}+\lambda_{4})\lambda_{6}+2C_{2}^{2}(\lambda_{2}+\lambda_{5})\lambda_{6}+(B_{1}-B_{2}+C_{2})^{2}\lambda_{6}^{2}.$ (56e)
@@ Appendix G Thermodynamic Jacobian calculations
$\dfrac{\partial\left(P,x_{p}-x_{\mu},\delta\mu_{2}\right)}{\partial\left(n_{B},x_{p},x_{\mu}\right)}\bigg{/}\dfrac{\partial\left(x_{p}-x_{\mu},\delta\mu_{2}\right)}{\partial\left(x_{p},x_{\mu}\right)}=\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{x_{p},x_{\mu}}-\left(\dfrac{\partial P}{\partial x_{p}}\dfrac{\partial\delta\mu_{2}}{\partial n_{B}}+\dfrac{\partial P}{\partial x_{\mu}}\dfrac{\partial\delta\mu_{2}}{\partial n_{B}}\right)\bigg{/}\left(\dfrac{\partial\delta\mu_{2}}{\partial x_{\mu}}+\dfrac{\partial\delta\mu_{2}}{\partial x_{p}}\right)$

### v3

@@ IV.2 Particle dynamics during an oscillation
The evolution of the particle fractions $x_{p}(t)$, $x_{\mu}(t)$, and $x_{\pi}(t)$ are plotted in Fig. 10 at different temperatures and densities, with and without pions in the EoS. These quantities are obtained from Eq. 12, where the real and imaginary parts of $\delta x_{i}$ are the solutions of the system of equations described after Eq. 27. Note that we plot $\delta x_{i}(t)$, which indicates how far the particle fraction of species $i$ is from its beta equilibrium value at the background density, not the concurrent value of the density $n_{B}(t)$.
In studying these plots, both the height and the phase of the sinusoidal variations in the particle fractions are informative. Our intuition for the behavior of $\delta x_{i}(t)$ comes from the $npe$ matter case. For example, Eq. 34 in the arXiv version of [14] indicates that in cold matter, where the beta equilibration is slow, $\delta x_{p}(t)\sim(\delta n_{B}/n_{B})\left[A/(|B|n_{B})\right](\gamma/\omega)\sin{(\omega t)}$, which [^8] Here, $A$ and $B$ correspond to $A_{1}$ and $B_{1}$ in this text (removing, of course, the designation of fixed muon and pion fractions as the expression in Ref. [14] applies to $npe$ matter). is out of phase with the density oscillation, but also vanishes in the slowly equilibrating limit (that is, when the Urca rate is too slow to chemically equilibrate the matter, the particle fraction does not change from the beta equilibrium value at the background density). In the opposite limit, where the beta equilibration is very fast, $\delta x_{p}(t)\sim(\delta n_{B}/n_{B})\left[A/(|B|n_{B})\right]\cos{(\omega t)}$. That is, the particle fraction oscillation is in phase with the density oscillation and, at the maximum in the $x_{p}(t)$ oscillation cycle, has a magnitude that is independent of the equilibration rate. The particle fraction just tracks the beta equilibrium curve $x_{i}^{\beta-eq.}[n_{B}(t)]$ throughout the density oscillation.
While the equilibration rate $\gamma$ is a strong function of $T$, it is not necessarily the case that the above description of $\delta x_{p}(t)$ at low and high $\gamma$ can be mapped to low and high temperature, as the susceptibility prefactor in $\delta x_{p}(t)$ depends on temperature too. For low temperature $npe$ matter, the susceptibilities are relatively temperature-independent. But for a system including thermal pions, the susceptibilites are not independent of temperature, as we have seen (c.f. Figs. 3 and 4).
Without pions (dashed lines in Fig. 10), we see the expected behavior. The fractions $x_{p}(t)$ and $x_{\mu}(t)$ deviate little from their beta equilibrium values at low temperature, and as temperature increases, they deviate further, as the equilibration is fast enough to force the particle fractions to trace the beta equilibrium curve $x_{i}^{\beta-eq.}[n_{B}(t)]$. The particle fraction oscillations also become more in-phase with the density oscillation with increasing temperature.
When pions are included in the EoS, the expected behavior (increasing deviation $\delta x_{i}(t)$ with increasing temperature, up to some maximum deviation) still occurs at $1n_{0}$, but at $1.6n_{0}$, the trend reverses for some of the particle species. Indeed, the presence of the pions (through their contribution to the susceptibilities) causes the particle fractions $x_{p}(t)$ and $x_{\pi}(t)$ to oscillate less dramatically as temperature increases, while the muon fraction $x_{\mu}(t)$ oscillation amplitude increases in the expected way with increasing temperature.
@@ V Conclusions
Our main finding is that thermal pions can significantly alter the density and temperature dependence of the bulk viscosity of hot dense matter. The pion population rapidly increases with density and temperature, and so the bulk viscosity deviates most strongly from the pionless case at higher densities and temperatures. The addition of pions to the EoS modifies the existing nuclear matter susceptibilities to a relatively small extent (Figs. 3 and 4) but also introduces new susceptibilities that involve the pion degree of freedom and therefore do not exist without pions. These new susceptibilities have a significant effect on the bulk viscosity. In addition, pion reaction rates have two different effects. The (essentially) infinitely fast strong interaction $n+n\rightarrow n+p+\pi^{-}$ does not produce a new resonant peak in the bulk viscosity in the temperature range studied, but does (as one can see from Fig. 5) modify the maximum value of the bulk viscosity and can move or even eliminate or introduce partial conformal points in the bulk viscosity at fixed density, $\zeta(T)$. The slower pion decays ($\pi^{-}\rightarrow\mu^{-}+\bar{\nu}_{\mu}$) in some conditions can overtake the existing nucleonic or leptonic equilibration processes and can dominate beta equilibration.
@@ Acknowledgements.
SPH thanks Mark Alford for collaboration in the early stages of the project, Albino Perego and Eleonora Loffredo for information about the conserved lepton fractions in neutron star mergers, and Alex Haber for comments on the draft of this paper. The work of SPH was supported by the U.S. Department of Energy grant DE-FG02-00ER41132 as well as the National Science Foundation grants PHY-1430152 (JINA Center for the Evolution of the Elements) and PHY 21-16686. BF is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under contracts DE-AC02-06CH11357, by the 2020 DOE Early Career Award number ANL PRJ1008597, by the NUCLEI SciDAC program, and Argonne LDRD awards. The work of SR was supported by the U.S. Department of Energy grant DE-FG02-00ER41132. SR also thanks the members of the N3AS Physics Frontier Center, funded by the NSF Grant No. PHY-2020275 for useful conversations.
@@ Appendix A Bulk viscosity of matter with net muon lepton fraction $Y_{L\mu}$
The particle fractions in matter with net conserved muon number are plotted in Fig. 11. When $Y_{Le}=Y_{L\mu}=0.05$, the electron and muon populations are roughly equal. At $n_{0}$, the proton fraction is 10% instead of around 5% (in the $Y_{L\mu}=0$ case). When pions are added to the system, their population exponentially rises with temperature, becoming appreciable at $T\approx 20\text{ MeV}$. When the pion population becomes appreciable, the proton fraction is raised and the lepton fractions are slightly lowered. At $1.6n_{0}$, the pions content becomes appreciable at very low temperatures, just a couple MeV. Thus the proton fraction is larger than without pions for the basically the entire temperature range plotted.
@@
Figure 12: Susceptibilities of the $A$ type from the EoS described in Sec. II plotted in merger conditions (top panels) and in supernovae conditions (bottom panels), but with equal conserved lepton fractions. The dashed lines correspond to an EoS without pions, while the solid lines indicate an EoS with pions. $A_{3}$ has no counterpart in a system without pions.
Figure 13: Susceptibilities of the $B$, $C$, and $D$ types from the EoS described in Sec. II plotted in merger conditions (top panels) and in supernovae conditions (bottom panels), but with equal conserved lepton fractions. The dashed lines correspond to an EoS without pions, while the solid lines indicate an EoS with pions. $D_{3}$ has no counterpart in a system without pions.
@@
The equilibration rates $\gamma_{i}$ are plotted in Fig. 14. In merger conditions, the rates are much faster when $Y_{L\mu}=0.05$ than when $Y_{L\mu}=0$. Adding pions to the EoS in these equal lepton fraction conditions modifies the beta equilibration rates slightly, especially as temperature rises. In supernovae conditions, the rates are hardly modified at all, because the pion content is very low.
@@ B.1 Direct Urca (with electrons)
The phase space integral reduces from 12 to 8 dimensions after integrating over the delta functions. Then a spherical coordinate system can be chosen in such a way to render three additional angular integrals trivial, leaving a 5-dimensional integral. Two more integrals can be done analytically, leaving a 3-dimensional integral to do numerically. This integral has been done with the full matrix element (Eq. 37) with $g_{A}=1$ in [4], with the non-relativistic matrix element 38 in [5], and in the neutrino-transparent case in [51, 50].
@@ B.2 Pion decay to electron
The rate of the process $\pi^{-}\rightarrow\mu^{-}+\bar{\nu}_{\mu}$ can be obtained by taking the rate integral 40 with matrix element 41 and replacing the electron mass and chemical potential with those of the muon, as well as replacing the electron neutrino chemical potential with the muon neutrino chemical potential.
@@ Appendix C Maxwell Relations
From this 1st law of thermodynamics, we can derive six Maxwell relations (if we only consider derivatives with respect to $n_{B}$ or $x_{i}$) [68]. The three that are relevant are written in Eqs. 23a, 23b, and 23c.
@@ Appendix F Bulk viscosity $\lambda_{3}\rightarrow\infty$
$+\lambda_{6}\bigg{[}A_{2}(B_{1}-B_{2})+A_{1}C_{2}-\dfrac{A_{3}B_{2}(B_{1}-B_{2}+C_{2})}{B_{2}+D_{3}}\bigg{]}^{2}\Big{\}}$
@@
$+2(B_{1}-B_{2})^{2}(\lambda_{1}+\lambda_{4})\lambda_{6}+2C_{2}^{2}(\lambda_{2}+\lambda_{5})\lambda_{6}+(B_{1}-B_{2}+C_{2})^{2}\lambda_{6}^{2}.$
@@ Appendix G Thermodynamic Jacobian calculations
$=\dfrac{\partial\left(P,x_{p}-x_{\mu},\delta\mu_{2}\right)}{\partial\left(n_{B},x_{p},x_{\mu}\right)}\bigg{/}\dfrac{\partial\left(x_{p}-x_{\mu},\delta\mu_{2}\right)}{\partial\left(x_{p},x_{\mu}\right)}=\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{x_{p},x_{\mu}}-\left(\dfrac{\partial P}{\partial x_{p}}\dfrac{\partial\delta\mu_{2}}{\partial n_{B}}+\dfrac{\partial P}{\partial x_{\mu}}\dfrac{\partial\delta\mu_{2}}{\partial n_{B}}\right)\bigg{/}\left(\dfrac{\partial\delta\mu_{2}}{\partial x_{\mu}}+\dfrac{\partial\delta\mu_{2}}{\partial x_{p}}\right)$